A car goes down a certain road at an average speed of 40 km/h and returns along the same road at an average speed of 60 km/h.  Calculate the average speed for the round trip.

Grade:12

3 Answers

Chetan Mandayam Nayakar
13 years ago

let d be the one way distance. average speed for round trip= 2d/((d/40)+(d/60))=2*40*60/(40+60)=48 km/h

vikas askiitian expert
13 years ago

let the distance between the two points be D

   time taken by car to reach the other point is t1=D/40

  time taken by car to come back to the initial point is t2=D/60

   total distance travelled by car is 2D

                  average speed = total distance /total time

                                        =2D/(D/40 + D/60)

                                         =48km/h

rajiv kumar
13 years ago

50 m/s.